# Achieving Optimal Sample Complexity for a Broader Class of Signals in Sparse Phase Retrieval

### v2

@@ Appendix A Basic tools and Lemmas
Denote the smallest $\epsilon$-net of $T$ as $N_{\epsilon}^{T}$, which means that $|N_{\epsilon}^{T}|=\mathcal{N}(T,\epsilon)$. Then, for any unit $\mathbf{z}_{1},\mathbf{z}_{2}\in T$, there exist $\mathbf{z}_{10}\in N_{\epsilon}^{T}$ and $\mathbf{z}_{20}\in N_{\epsilon}^{T}$ such that $\|\mathbf{z}_{1}-\mathbf{z}_{10}\|\leq\epsilon$ and $\|\mathbf{z}_{2}-\mathbf{z}_{20}\|\leq\epsilon$. It holds that
$\mathbf{z}_{1}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{2}$
== $\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{20}+\mathbf{z}_{1}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)(\mathbf{z}_{2}-\mathbf{z}_{20})$
$+(\mathbf{z}_{1}-\mathbf{z}_{10})^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{20}$
$\overset{(a)}{\leq}$ $\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{20}+2\epsilon\left\lVert\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right\rVert,$
where $(a)$ comes from the definition of $N_{\epsilon}^{T}$ and the matrix norm, together with the fact that $\|\mathbf{z}_{20}\|\leq 1$. We first prove the following two inequalities
$\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\right)\mathbf{z}_{20}\leq\frac{\eta}{4}\times\frac{1}{g(k,s(p))},$ (A.10)
$\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}\right)\mathbf{z}_{20}\leq\frac{\eta}{4}\times\frac{1}{g(k,s(p))}.$ (A.11)
$\dagger$} Proof for (A.10)
Simple calculation gives that
$\mathbf{z}_{10}^{*}\tilde{\mathbf{Z}}_{S}\mathbf{z}_{20}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{\mathbf{y}_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20},$
which is the sum of independent sub-exponential random variables with constant $\psi_{1}$ norms, since $\|\mathbf{z}_{10}\|,\|\mathbf{z}_{20}\|\leq 1$. Employing Bernstein’s inequality yields
$\mathbb{P}\left(\left\lvert\mathbf{z}_{10}^{*}\left(\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\right)\mathbf{z}_{20}\right\rvert>\frac{\eta}{4}\times\frac{1}{g(k,s(p))}\right)$
$\leq$~ $2\exp\left(-C^{\prime}_{\eta}m\min\left\{\frac{\eta^{2}}{16K^{2}g^{2}(k,s(p))},\frac{\eta}{4Kg(k,s(p))}\right\}\right)$
$\doteq$~ $p_{1},$
where the sub-exponential norm satisfies
$K=\max_{i\in[m]}\left\rVert\left(\frac{1}{2}-\exp\left(-\frac{\mathbf{y}_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\left\lvert(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{10}\right\rvert\left\lvert(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert\right\rVert_{\psi_{1}}$
and $c_{\eta}$ is the numerical constant.
Take the union bound for all $\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}$, we have
$~\mathbb{P}\left(\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\right)\mathbf{z}_{20}\leq\frac{\eta}{4}\times\frac{1}{g(k,s(p))}\right)$
$\geq$ $~1-\mathcal{N}^{2}(T,\epsilon)\times p_{1}.$
Furthermore, take the union bound for all $S\subseteq[n]$ satisfying $|S|\leq q$, the probability will exceed $1-{n\choose q}\mathcal{N}^{2}(T,\epsilon)p_{1}$.
Therefore, when
$m\geq C_{\eta}^{\prime}g^{2}(k,s(p))(\log\mathcal{N}(T,\epsilon)+q\log(n/q)),$
(A.10) holds with probability exceeding $1-2\exp\big(c^{\prime}_{\eta}m\big)$.
$\dagger$} Proof for (A.11)
Direct calculation yields
$\left\rvert\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}\right)\mathbf{z}_{20}\right\rvert$
@@
== $\left\lvert\frac{1}{m}\sum_{i=1}^{m}\left(\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)-\exp\left(-\frac{y_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\mathbf{z}_{10}^{*}(\mathbf{a}_{i}\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert$ (A.14)
$\overset{(a)}{=}$ $\max_{i\in[m]}\exp\left(-\frac{y_{i}^{2}}{\xi}\right)\hskip-0.7113pt\frac{y_{i}^{2}}{\xi^{2}}\hskip-0.7113pt\big|\lambda^{2}-\|\mathbf{x}\|^{2}\big|\hskip-0.7113pt\hskip-0.7113pt\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i}\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert$ (A.17)
$\overset{(b)}{\leq}$ $~\frac{1}{\xi}\big|\lambda^{2}-\|\mathbf{x}\|^{2}\big|\times\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert$, (A.20)
where $(a)$ comes from Lagrange’s mean value formula with $F(t)=\exp\big(\frac{y_{i}^{2}}{t}\big)$ and $\xi\in\left[\min\{\lambda^{2},\|\mathbf{x}\|^{2}\},\max\{\lambda^{2},\|\mathbf{x}\|^{2}\}\right]$, while $(b)$ employs the numerical inequality $xe^{-x}\leq 1,\ x\geq 0$.
For the second term, since $\mathbf{a}_{i}$’s are complex standard Gaussian random vectors and $\|\mathbf{z}_{10}\|,\|\mathbf{z}_{20}\|\leq 1$, it can be derived that for every $i\in[m]$, $\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}$ is sub-exponential random variable with constant $\psi_{1}$ norm and expectation being $(\mathbf{z}_{10}^{*})_{S}(\mathbf{z}_{20})_{S}$. Therefore, follow the same concentration analysis as in the proof of (A.10), we conclude that the inequality
$\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}-(\mathbf{z}_{10}^{*})_{S}(\mathbf{z}_{20})_{S}\right\rvert\leq\frac{\eta}{12}\times\frac{1}{g(k,s(p))}$ (A.21)
holds for all $\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}$ and $S\subseteq[n]$ satisfying $|S|\leq q$ with probability exceeding $1-2\exp(c^{\prime\prime}_{\eta}m)$ when
$m\geq C^{\prime\prime}_{\eta}g^{2}(k,s(p))(\log\mathcal{N}(T,\epsilon)+q\log(n/q)).$
Note that
$\left\lvert(\mathbf{z}_{10}^{*})_{S}(\mathbf{z}_{20})_{S}\right\rvert\leq\sqrt{\|\mathbf{z}_{10}\|^{2}\|\mathbf{z}_{20}\|^{2}}\leq 1.$
The inequality (A.21) can be transformed to
$\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert\leq 1+\frac{\eta}{12}\times\frac{1}{g(k,s(p))}.$ (A.22)
For the first term, by employing Lemma 3 with with $\kappa=g(k,s(p))$ and $\beta=\eta/12$, we conclude that
$\frac{1}{\xi}\left\lvert\lambda^{2}-\|\mathbf{x}\|^{2}\right\rvert$ $\leq\frac{1}{\|\mathbf{x}\|^{2}-\left\lvert\lambda^{2}-\|\mathbf{x}\|^{2}\right\rvert}\left\lvert\lambda^{2}-\|\mathbf{x}\|^{2}\right\rvert$
$\leq\frac{1}{\left(1-\frac{\eta}{12g(k,s(p))}\right)\|\mathbf{x}\|^{2}}\times\frac{\eta\|\mathbf{x}\|^{2}}{12g(k,s(p))}$ (A.23)
@@
holds true with probability exceeding $1-e^{-cm}$ as soon as
$m\geq C^{\prime\prime}_{\eta}g^{2}(k,s(p)).$
Finally, taking (A.22) and (A.23) into (A.20) yields
[…]
$\overset{(a)}{\leq}$ $~3\times\frac{\eta}{12}\times\frac{1}{g(k,s(p))}=\frac{\eta}{4}\times\frac{1}{g(k,s(p))},$
where $(a)$ comes from the property $\frac{1+x}{1-x}\leq 3$ with $x\leq\frac{1}{2}$ and the fact that $\frac{\eta}{12}\times\frac{1}{g(k,s(p))}\leq\frac{1}{2}$. Here, we complete the proof of two sub-conclusions. Finally, combining (A.10) and (A.11) with $T=\{\mathbf{u}|\operatorname{supp}(\mathbf{u})=S,\|\mathbf{u}\|\leq 1\}$ yields (A.9), which in turn proves our conclusion. ∎
Remark 4.
As indicated above, the term $q\log(n/q)$ arises from the union bound for $S\in[n],|S|=q$. For fixed set $S$, the sample complexity reduces to
$m\geq C_{\eta}g^{2}(k,s(p))\log\mathcal{N}(T,\epsilon).$ (A.24)
@@ A-A Concentration analysis for the sum of the diagonal entries
For all $\mathcal{S}\subseteq[n]$ with $|\mathcal{S}|=p$ and constant $0<\eta<1$, the following concentration inequality
$\left\lvert\frac{1}{p}\sum_{j\in\mathcal{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert<\eta\times\frac{1}{ps(p)}$ (A.25)
holds true with probability exceeding $1-2e^{-c_{\eta}m}$ given $m\geq C_{\eta}p^{2}s^{2}(p)\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constants related to $\eta$.
@@
We can see from (A.3) that
$\mathbb{E}[\tilde{Z}_{j}]=\left\{\begin{array}[]{cc}0,&j\notin{\text{supp}(\mathbf{x})},\\ \frac{|x_{j}|^{2}}{4\|\mathbf{x}\|^{2}},&j\in{\text{supp}(\mathbf{x})}.\end{array}\right$. (A.26)
Therefore, the left term of (A.25) can be expressed as
$\frac{1}{p}\sum_{j\in\mathcal{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)=\frac{1}{p}\sum_{j\in\mathcal{S}}\left(\mathbf{Z}_{\{j\}}\hskip-0.7113pt-\hskip-0.7113pt\frac{\mathbf{x}_{\{j\}}\mathbf{x}_{\{j\}}^{*}}{4\|\mathbf{x}\|^{2}}\right).$
For all $j\in[n]$, employing Lemma 4 with $S=\{j\}$, the fixed $T=\{\mathbf{1},-\mathbf{1}\}$, and $g(k,s(p))=ps(p)$, we can derive that
$\sup_{j\in[n]}\mathbf{1}^{*}\left(\mathbf{Z}_{\{j\}}-\frac{\mathbf{x}_{\{j\}}\mathbf{x}_{\{j\}}^{*}}{4\|\mathbf{x}\|^{2}}\right)\mathbf{1}\leq\eta\times\frac{1}{ps(p)},$
$\sup_{j\in[n]}\mathbf{1}^{*}\left(\mathbf{Z}_{\{j\}}-\frac{\mathbf{x}_{\{j\}}\mathbf{x}_{\{j\}}^{*}}{4\|\mathbf{x}\|^{2}}\right)(-\mathbf{1})\leq\eta\times\frac{1}{ps(p)}.$
holds with probability exceeding $1-2e^{-c_{\eta}m}$ and the corresponding sample complexity becomes
$m\geq C_{\eta}p^{2}s^{2}(p)\log n,$
which in turn proves (A.25) since $\mathcal{S}\subseteq[n]$.
@@
holds with probability exceeding $1-2\exp(-c_{\eta}m)$ given $m\geq C_{\eta}\frac{ks(p)}{\gamma^{2}}\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constant related to $\eta$.
@@
Note that the expression of $f_{l}$ can be transformed as
$f_{l}=\mathbf{e}_{l}^{*}\mathbf{Z}_{S^{0}}\mathbf{e}^{0},$
where $\mathbf{e}_{l}\in{\mathbb{C}}^{n}$ is the vector that sets its $l$-th entry as 1 while others as 0, $S^{0}$ is a fixed set derived by Step 1 of gESP and $\mathbf{e}^{0}$ is the principle eigenvector of $\mathbf{Z}_{S^{0}}$. Employing Lemma 4 with fixed $S=S^{0}$, fixed $T=\{\mathbf{e}_{1},\mathbf{e}_{2},\cdots,\mathbf{e}_{n},\mathbf{e}^{0}\}$ and $g(k,s(p))=\frac{2}{\gamma}\sqrt{ks(p)}$, we derive that
$\max_{l\in[n]}\left\lvert f_{l}-\mathbb{E}(\tilde{f}_{l})\right\rvert<\eta\times\frac{\gamma}{2\sqrt{ks(p)}}$
holds true with probability exceeding $1-2\exp(-c_{\eta}m)$ with the sample complexity satisfying $m\geq C_{\eta}\frac{ks(p)}{\gamma^{2}}\log n$.

### v3

@@ Appendix A Basic Tools and Lemmas
Note that the scalar coefficients $\xi_{i}=\frac{1}{2}-\exp(-y_{i}^{2}/\|\mathbf{x}\|^{2})$ are strictly bounded, since $|\xi_{i}|\leq 1/2$. By applying Lemma 3, we obtain the desired result. ∎
Next, we bound the operator norm of the difference matrix $\mathbf{Z}-\tilde{\mathbf{Z}}$.
Lemma 5.
Let $\mathbf{Z}$ and $\tilde{\mathbf{Z}}$ be defined as in (A.1) and (A.2). For all $S\subseteq[n]$ with $|S|=s$, the following bound holds for the operator norm with probability exceeding $1-2\exp(-cs\log(n/s))$:
$\|\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}\|\leq C\sqrt{\frac{\log n}{m}}.$ (A.10)
Proof.
Write the difference matrix $\Delta\mathbf{Z}_{S}=\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}$:
$\Delta\mathbf{Z}_{S}=\frac{1}{m}\sum_{i=1}^{m}\left[\exp\left(-\frac{y_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)\right](\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}.$ (A.11)
Define the scalar function $g(t)=\exp(-y_{i}^{2}/t)$. By the Lagrange Mean Value Theorem, there exists a $\xi_{i}$ between $\min(\lambda^{2},\|\mathbf{x}\|^{2})$ and $\max(\lambda^{2},\|\mathbf{x}\|^{2})$ such that
$g(\|\mathbf{x}\|^{2})-g(\lambda^{2})=g^{\prime}(\xi_{i})(\|\mathbf{x}\|^{2}-\lambda^{2}).$
The derivative is $g^{\prime}(t)=\frac{y_{i}^{2}}{t^{2}}\exp(-y_{i}^{2}/t)$. Thus,
$\exp\left(-\frac{y_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)=\underbrace{\frac{y_{i}^{2}}{\xi_{i}^{2}}\exp\left(-\frac{y_{i}^{2}}{\xi_{i}}\right)}_{\delta_{i}\geq 0}(\|\mathbf{x}\|^{2}-\lambda^{2}).$
Substituting this back into the matrix sum yields
$\Delta\mathbf{Z}_{S}=(\|\mathbf{x}\|^{2}-\lambda^{2})\cdot\left(\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right).$
Now, we apply the operator norm $\|\cdot\|$. Since $|\lambda^{2}-\|\mathbf{x}\|^{2}|$ is a global scalar, we can pull it out and obtain
$\|\Delta\mathbf{Z}_{S}\|=|\|\mathbf{x}\|^{2}-\lambda^{2}|\cdot\left\|\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|.$ (A.12)
First, we bound the scalar coefficient $\delta_{i}$ uniformly. Consider the function $h(u)=ue^{-u}$ for $u\geq 0$. Its maximum value is $e^{-1}$ at $u=1$.We can rewrite $\delta_{i}$ as
$\delta_{i}=\frac{1}{\xi_{i}}\left(\frac{y_{i}^{2}}{\xi_{i}}\exp\left(-\frac{y_{i}^{2}}{\xi_{i}}\right)\right)\leq\frac{1}{\xi_{i}}\cdot e^{-1}.$
Note that for each $i$, $\xi_{i}$ is between $\min(\lambda^{2},\|\mathbf{x}\|^{2})$ and $\max(\lambda^{2},\|\mathbf{x}\|^{2})$. From Lemma 2, denote $\epsilon=C\sqrt{\frac{\log n}{m}}$, and we have $\xi_{i}\geq(1-\epsilon)\|\mathbf{x}\|^{2}$. Thus,
$0\leq\delta_{i}\leq\frac{1}{e(1-\epsilon)\|\mathbf{x}\|^{2}}\doteq\frac{C_{1}}{\|\mathbf{x}\|^{2}}.$ (A.13)
Second, we bound $\left\|\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|$ via the PSD property. Since $(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}$ are Positive Semidefinite (PSD) matrices and coefficients $\delta_{i}$ are non-negative, it holds that
$\left\|\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|\leq\left(\max_{i}\delta_{i}\right)\left\|\frac{1}{m}\sum_{i=1}^{m}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|.$
Using the bound (A.13) for $\delta_{i}$ and applying Lemma 3 with $\xi_{i}\equiv 1$, we have
$\left\|\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|\leq\frac{C_{1}}{\|\mathbf{x}\|^{2}}\left(1+C_{2}\sqrt{\frac{s\log(n/s)}{m}}\right).$
Finally, substituting this back into equation (A.12) yields
@@
$\|\Delta\mathbf{Z}_{S}\|\leq|\|\mathbf{x}\|^{2}-\lambda^{2}|\cdot\frac{C_{1}}{\|\mathbf{x}\|^{2}}\left(1+C_{2}\sqrt{\frac{s\log(n/s)}{m}}\right).$
Applying Lemma 2, we have
$\|\Delta\mathbf{Z}_{S}\|\leq C_{1}C_{3}\sqrt{\frac{\log n}{m}}\cdot\left(1+C_{2}\sqrt{\frac{s\log(n/s)}{m}}\right).$
Choosing $m\geq C_{4}s\log(n/s)$ with sufficiently large $C_{4}$ allows the union bound to control the failure probability by $2\exp(-cs\log(n/s))$. This completes the proof. ∎
As a consequence, via triangle inequality, we have the following result.
Corollary 4.
For all $S\subseteq[n]$ with $|S|=s$, the following concentration inequality
$\|\mathbf{Z}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\|\leq C\sqrt{\frac{s\log(n/s)}{m}}$ (A.14)
@@
holds true with probability exceeding $1-2e^{-cs\log(n/s)}$ given $m\geq Cs\log(n/s)$, where $c$ and $C$ are numerical constants.
@@ A-A Concentration analysis for the sum of the diagonal entries
For all $S\subseteq[n]$ with $|S|=p$ and constant $0<\eta<1$, the following concentration inequality
$\left\lvert\frac{1}{p}\sum_{j\in S}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert<\eta\times\frac{1}{ps(p)}$ (A.15)
holds true with probability exceeding $1-n^{-c_{\eta}}$ given $m\geq C_{\eta}p^{2}s^{2}(p)\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constants related to $\eta$.
@@
The left term of (A.15) can be expressed as
$\frac{1}{p}\sum_{j\in S}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)$ $=\frac{1}{p}\sum_{j\in S}\left(\mathbf{Z}_{\{j\}}-\mathbb{E}[\tilde{\mathbf{Z}}_{\{j\}}]\right)$
$\leq\max_{j\in S}\left\lvert\mathbf{Z}_{\{j\}}-\mathbb{E}[\tilde{\mathbf{Z}}_{\{j\}}]\right\rvert$.
For all $j\in[n]$, employing Corollary 4 with $S=\{j\}$ (and thus $|S|=1$), we have
$\max_{j\in S}\left\lvert\mathbf{Z}_{\{j\}}-\mathbb{E}[\tilde{\mathbf{Z}}_{\{j\}}]\right\rvert\leq C\sqrt{\frac{\log n}{m}}.$
Taking $m\geq C\eta^{-2}p^{2}s^{2}(p)\log n$ yields the desired result.
@@
holds with probability exceeding $1-n^{-c_{\eta}}$ given $m\geq C_{\eta}{\gamma^{-2}}ks(p)\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constant depending on $\eta$.
@@
The goal is to bound the element-wise error $\max_{l}|f_{l}-\mathbb{E}[\tilde{f}_{l}]|$. We rely on the independence between the measurement vectors $\mathbf{a}_{i}$ in the current stage and the vector $\mathbf{e}^{0}$. We decompose the error into two terms as follows:
$|f_{l}-\mathbb{E}[\tilde{f}_{l}]|$ $\leq\underbrace{|f_{l}-\tilde{f}_{l}|}_{\text{bias term}}+\underbrace{|\tilde{f}_{l}-\mathbb{E}[\tilde{f}_{l}]|}_{\text{fluctuation term}}$
$=|\mathbf{e}_{l}^{*}(\mathbf{Z}-\tilde{\mathbf{Z}})\mathbf{e}^{0}|+|\mathbf{e}_{l}^{*}(\tilde{\mathbf{Z}}-\mathbb{E}[\tilde{\mathbf{Z}}])\mathbf{e}^{0}|.$ (A.17)
On one hand, the bias term is controlled by the operator norm of the difference matrix $\Delta\mathbf{Z}=\mathbf{Z}-\tilde{\mathbf{Z}}$. Note that $\mathbf{e}^{0}$ is supported on $S^{0}$, and denote $U=S^{0}\cup\{l\}$ of size $|U|=p+1\leq 2k$. Then we have
$|\mathbf{e}_{l}^{*}(\mathbf{Z}-\tilde{\mathbf{Z}})\mathbf{e}^{0}|$ $\makebox[61.83356pt][c]{$\leq$}\|(\mathbf{Z}-\tilde{\mathbf{Z}})_{U}\|\|\mathbf{e}_{l}\|_{2}\|\mathbf{e}^{0}\|_{2}$ (A.20)
$\makebox[61.83356pt][c]{$\leq$}\|(\mathbf{Z}-\tilde{\mathbf{Z}})_{U}\|$ (A.23)
$\makebox[61.83356pt][c]{$\overset{(a)}{\leq}$}C\sqrt{\frac{\log n}{m}},$ (A.26)
where (a) follows from Lemma 5.
On the other hand, conditioned on $S^{0}$ (and thus $\mathbf{e}^{0}$), we analyze the fluctuation term $|\tilde{f}_{l}-\mathbb{E}[\tilde{f}_{l}]|$. Recall that $\tilde{\mathbf{f}}=\tilde{\mathbf{Z}}\mathbf{e}^{0}$. We can expand the $l$-th entry as a sum of $m$ independent random variables:
$\tilde{f}_{l}=\mathbf{e}_{l}^{*}\tilde{\mathbf{Z}}\mathbf{e}^{0}=\frac{1}{m}\sum_{i=1}^{m}\kappa_{i}(\mathbf{a}_{i}^{*}\mathbf{e}_{l})^{*}(\mathbf{a}_{i}^{*}\mathbf{e}^{0}),$ (A.27)
where $\kappa_{i}=\frac{1}{2}-\exp(-y_{i}^{2}/\|\mathbf{x}\|^{2})$. Since we assume independence between the current samples $\{\mathbf{a}_{i}\}$ and the initialization $\mathbf{e}^{0}$ (data splitting), the terms in the summation are independent conditioned on $\mathbf{e}^{0}$. Let $u_{i}=\mathbf{a}_{i}^{*}\mathbf{e}_{l}$ and $v_{i}=\mathbf{a}_{i}^{*}\mathbf{e}^{0}$. Both $u_{i}$ and $v_{i}$ are standard complex Gaussian variables (potentially correlated if $l\in\text{supp}(\mathbf{e}^{0})$). The $i$-th summand is $X_{i}=\kappa_{i}\bar{u}_{i}v_{i}$. Note that the weights are bounded as $|\kappa_{i}|\leq 1/2$. The product of two Gaussian variables $\bar{u}_{i}v_{i}$ follows a sub-exponential distribution with bounded Orlicz norm $\|\bar{u}_{i}v_{i}\|_{\psi_{1}}\leq C\|u_{i}\|_{\psi_{2}}\|v_{i}\|_{\psi_{2}}\leq C^{\prime}$. Consequently, the weighted variable $X_{i}$ is sub-exponential with
$\|X_{i}\|_{\psi_{1}}\leq|\kappa_{i}|\|\bar{u}_{i}v_{i}\|_{\psi_{1}}\leq C^{\prime\prime}.$ (A.28)
Applying the scalar Bernstein inequality [23, Theorem 2.8.1] directly, we have
$\mathbb{P}(|\tilde{f}_{l}-\mathbb{E}[\tilde{f}_{l}]|\leq t)\geq 1-2\exp\left(-cm\min\{t,t^{2}\}\right).$
Setting $t=\sqrt{\frac{\log n}{m}}$ and taking the union bound over all $l\in[n]$, we get
$\max_{l}|\tilde{f}_{l}-\mathbb{E}[\tilde{f}_{l}]|\leq C^{\prime}\sqrt{\frac{\log n}{m}}.$ (A.29)
Combining (A.26) and (A.29), we have
$|f_{l}-\mathbb{E}[\tilde{f}_{l}]|\leq C\sqrt{\frac{\log n}{m}}.$ (A.30)
Therefore, when $m\geq C\eta^{-2}\gamma^{-2}ks(p)\log n$, the desired result holds with probability exceeding $1-n^{-c_{\eta}}$. ∎
A-C Proof for Example 1
For sufficiently large $k$, we have $k^{-\delta}\gg k^{-1/2}$ (since $\delta<1/2$), so the ordering in (49) is consistent with $x_{(1)}\geq x_{(2)}\geq\cdots\geq x_{(k)}$.
First, we have
$\|\mathbf{x}\|^{2}=k^{-2\delta}+(m-1)\cdot\frac{1}{4}k^{-2\delta}+(k-m)\cdot k^{-1}=\Theta(1),$ (A.31)
because $(m-1)k^{-2\delta}=\Theta\!\big(k^{1/2-2\delta}\big)=o(1)$ for $\delta>1/4$, and $(k-m)k^{-1}=1-\Theta(k^{-1/2})$.
We now compare the orders of magnitude for the quantities of interest:
  • For $p=1$,
    $s(1)=\frac{\|\mathbf{x}\|^{2}}{|x_{(1)}|^{2}}=\Theta(1)\cdot k^{2\delta}=\Theta\!\big(k^{2\delta}\big).$ (A.32)
  • For $m=\lceil\sqrt{k}\rceil$,
    $\sum_{j=1}^{m}|x_{(j)}|^{2}=k^{-2\delta}+(m-1)\cdot\frac{1}{4}k^{-2\delta}=\Theta\!\big(k^{1/2-2\delta}\big).$ (A.33)
    Thus we have
    $s^{2}(m)=\Theta\!\big(k^{4\delta-1}\big).$ (A.34)
Consequently, for all $\delta\in(1/4,1/2)$, we have
$\frac{s^{2}(m)}{s(1)}=\Theta\!\big(k^{2\delta-1}\big)\to 0,$ (A.35)
and hence $\Theta(s^{2}(\lceil\sqrt{k}\rceil))<\Theta(s(1))$.